) It is approximately 480 miles from Los Angeles, California, to San Francisco, California. Allowing for various traffic conditions, a driver can average approximately 60 miles per hour.How far have you traveled after 3 hours?How far have you traveled after 4 hours?How far have you traveled after t hours (i.e., write a linear function that expresses the distance traveled, d, as a function of time, t).How far will you HAVE LEFT to travel to reach San Francisco after you have traveled 3 hours?How far will you HAVE LEFT to travel to reach San Francisco after you have traveled 4 hours?How far will you HAVE LEFT to travel to reach San Francisco after you have traveled t hours (i.e., write a linear function that expresses the distance to be traveled to reach San Francisco, s, as a function of time, t).
Answer by venugopalramana(3286)

It is approximately 480 miles from Los Angeles, California, to San Francisco, California. Allowing for various traffic conditions, a driver can average approximately 60 miles per hour.How far have you traveled after 3 hours?
DISTANCE TRAVELED = D = SPEED *TIME
DISTANCE TRAVELED IN 3 HRS = 60*3=180 MILES
How far have you traveled after 4 hours?
DISTANCE TRAVELED IN 4 HRS = 60*4=240 MILES
How far have you traveled after t hours (i.e., write a linear function that expresses the distance traveled, d, as a function of time, t).
DISTANCE TRAVELED IN T HRS....D = S*T MILES
How far will you HAVE LEFT to travel to reach San Francisco after you have traveled 3 hours?
TOTAL DISTANCE = 480 MILES
REMAINING DISTANCE = D' =480-D
DISTANCE TRAVELED IN 3 HRS = 60*3=180 MILES
REMAINING DISTANCE AFTER 3 HRS TRAVEL =480-180=300 MILES
How far will you HAVE LEFT to travel to reach San Francisco after you have traveled 4 hours?
REMAINING DISTANCE AFTER 4 HRS TRAVEL =480-240=240 MILES
How far will you HAVE LEFT to travel to reach San Francisco after you have traveled t hours (i.e., write a linear function that expresses the distance to be traveled to reach San Francisco, s, as a function of time, t).
REMAINING DISTANCE AFTER T HRS TRAVEL= D' =(480-ST) MILES